2) Contribution Margin

  1. DATA
    1. Fixed Costs
      1. Fixed Factory Overhead = $50,000
      2. Fixed Selling overhead = $10,000
    2. Variable Costs
      1. Variable Manufacturing costs = $10
      2. Variable selling cost per unit = $4
    3. Cost Per Unit = $35
  2. Calculate the CM
  3. Calculate the CM %

Note: the selling price is not given

How can we find the selling price by giving info
Answer #1

CM = price-variable cost = 35-4-10=21

CM% = CM/price = 21/35= 60%

media%2F8b0%2F8b03c98b-292d-45a6-a00a-be